Find three consecutive even integers such that the third is 16 less than twice the second
Answer by checkley77(12569)

Let x, x+2 & x+4 be the 3 integers.
x+4=2(x+2)-16
x+4=2x+4-16
x-2x=4-16-4
-x=-16
x=16 is for the first integer.
16+2=18 is the second integer.
16+4=20 is the third integer.
Proof:
20=2*18-16
20=36-16
20=20